OpenStudy (anonymous):

use the properties of logarithms to write the expression, log\[\sqrt[5]{3y/x}\] as a sum, difference, or multiples of logarithms of x, y, or constants. assume that x>0 and y>0

OpenStudy (anonymous):

its log of the rest of the stuff i wrote, it wouldnt go together

OpenStudy (anonymous):

1/5 {(log 3+log y)-log x}

OpenStudy (anonymous):

A=(3y/x)^(1/5) lnA=1/5[ln(3y/x)] lnA=1/5[ln3+lny-lnx] A=e^{1/5[ln3+lny-lnx]}